Suppose you have the following information on Islandland's production possiblity.

  • Islandland has 10000 labor hours available each year.
  • Producing cars requires 500 hours.
  • Producing tablets requires 50 hours.

Part A

Use this information to complete the table below. For Bundle A assume Islandland spends all labor hours producing cars. For Bundle B assumes Islandland spends 75% of labor hours producing cars and the remaining 25% on tablets. For Bundle C assume Islandland spends 25% of labor hours producing cars and the remaining 75% on tablets. For Bundle D assume Islandland spends 100% of labor hours producing tablets.

Round any answers 1 decimal place (example: 35.5).

Bundles A B C D
Cars
Tablets

Part B

Suppose a new technology allows for a reduction in the number of labor hours. Under the new technology it now requires only 25 hours to produce a tablet. How does this change the Production Schedule (assume the same percentages of labor hours for each bundle)?

Round any answers 1 decimal place (example: 35.5).

Bundles New Technology A B C D
Cars
Tablets
Answer #1

(Part A)

Bundles A B C D
Cars 10000/500 = 20.0 (10000 x 75%)/500 = 15.0 (10000 x 25%)/500 = 5.0 0.0
Tablets 0.0 (10000 x 25%)/50 = 50.0 (10000 x 75%)/50 = 150.0 10000/50 = 200.0

(Part B)

Bundles A B C D
Cars 10000/500 = 20.0 (10000 x 75%)/500 = 15.0 (10000 x 25%)/500 = 5.0 0.0
Tablets 0.0 (10000 x 25%)/25 = 100.0 (10000 x 75%)/25 = 300.0 10000/25 = 400.0